The number of tickets Cynthia sold is 12 less the number Holly sold. Together they sold 114. How many tickets did Holly sell?
Answer by TimothyLamb(4379)

---
x = Holly number of tickets
y = Cynthia number of tickets
---
y = x - 12
x + y = 114
---
put the system of linear equations into standard form:
---
x - y = 12
x + y = 114
---
copy and paste the above linear system in standard form into this matrix-method solver:
https://sooeet.com/math/system-of-linear-equations-solver.php
---
solution:
x = Holly number of tickets = 63
y = Cynthia number of tickets = 51
---
system matrix
1 -1
1 1
---
inverse of system matrix
0.5 0.5
-0.5 0.5
---
determinant of system matrix = 2
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Solve systems of linear equations up to 6-equations 6-variables:
https://sooeet.com/math/system-of-linear-equations-solver.php